OpenStudy (anonymous):

The area of a rectangle is 3,246 m^2. Find the dimension of the rectangle that gives the smallest perimeter for the rectangle

OpenStudy (anonymous):

\[Perimeter = 2*lengthOfsideA + 2*lengthOfsideB\] that is your constraint

OpenStudy (anonymous):

you have \[A=3246\] which is \[A=lengthOfsideA*lengthOfsideB\]

OpenStudy (anonymous):

a.b = 3246 2(a + b) should be min.

OpenStudy (anonymous):

Use the equation for Area to find a value for one of the sides in terms of the other e.g. \[lengthOfsideA = \frac{A}{lengthOfSideB}\]

OpenStudy (anonymous):

\[P=2(A+B)\] as Ishaan suggests and plug in the relation you discovered using the Area forumla, and then find the minimum of the function

OpenStudy (anonymous):

What are you saying?

OpenStudy (anonymous):

You do not know the lengths of either side of the rectangle

OpenStudy (anonymous):

You do know the area, and you do know the formula to get area

OpenStudy (anonymous):

you also know the formula for the perimeter of a rectangle

OpenStudy (anonymous):

ughhh. I dont get it

OpenStudy (anonymous):

a+b = 1623

OpenStudy (anonymous):

3246 is the area: area is length of side A * length of side B in the rectangle

OpenStudy (anonymous):

so \[3246 = A x B\]

OpenStudy (anonymous):

\[3246=A*B\] and you have \[Perimeter = 2*A + 2*B\]

OpenStudy (anonymous):

now, you want to use these two relationships

OpenStudy (anonymous):

so determine the length of A with regard to the area by dividing the area by the length of B \[Area=3246=A*B\] now \[A=\frac{3246}{B}\]

OpenStudy (anonymous):

OK, great, so you have \[A=\frac{3246}{B}\] and \[Perimeter = 2* A + 2*B\]

OpenStudy (anonymous):

now you can substitute for A in the Perimeter equation because you found that A is \[\frac{3246}{B}\]

OpenStudy (anonymous):

\[Perimeter = 2* (\frac{3246}{B}) + 2*B\]

OpenStudy (anonymous):

now, we can convert this into a quadratic polynomial like \[a*x^2 + b*x + c=0\] with a little manipulation

OpenStudy (anonymous):

first

OpenStudy (anonymous):

move Perimeter to the right side \[0 = 2*(\frac{3246}{B}) +2B - Perimeter\]then, multiply the entire thing by \[B\] because the B in the denominator \[\frac{3246}{B}\] is nasty and harder to deal with than this \[B*[2*(\frac{3246}{B})+ 2B - Perimeter] = \] is now \[2*(3246)+2B^2 - Perimeter*B=0\]

OpenStudy (anonymous):

rearrange this to be \[2B^2 - Perimeter*B + 2*3246=0\]

OpenStudy (anonymous):

omg. what was that? What is this called this is impossible

OpenStudy (anonymous):

whats this process called?

OpenStudy (anonymous):

this is just some substitution and algebra

OpenStudy (anonymous):

Okay. well.. I feel like i took two steps forward and five steps back

OpenStudy (anonymous):

check this out: if you have a rectangle with Area of 3246, if side A is 1 then side B would have to be 3246, that is super long and skinny rectangle

OpenStudy (anonymous):

the perimeter on that is \[2*3246+2*1\]

OpenStudy (anonymous):

that is huge

OpenStudy (anonymous):

so imagine you are manipulating the sides of the rectangle, like when you can draw a shape in photoshop and move your cursor around to change the dimensions

OpenStudy (anonymous):

you made a super long and skinny rectangle that has a huge perimeters, 3246 by 1, so you should probably squash it a bit, make it fatter

OpenStudy (anonymous):

is this better for conceptualizing what is going on?

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

go on..

OpenStudy (anonymous):

right, so you have the lengths of your rectangle

OpenStudy (anonymous):

A and B

OpenStudy (anonymous):

Perimeter of a rectangle is \[P=2*A + 2*B\] while \[Area=A*B\]

OpenStudy (anonymous):

so, if you are squashing the super long rectangle such that the area still 3246 but the lengths of the sides are changing, ultimately, what shape do you think will yield the shortest sides?

OpenStudy (anonymous):

I already got the answer. 56.97 by 56.97. Thanks